# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
In an era of rapid technological advancement, ensuring data security is of paramount importance. In this section, we propose a framework designed to address the security challenges in containerized and multi-cloud environment. This proposed framework, composed of seven categories of security challenges and solutions illustrated in Table 11, is derived through a comprehensive review of existing literature. It covers a broad spectrum of categories, each capturing specific security challenges and their corresponding solutions. Each category is briefly described below:
Data Security in container-based applications category covers a wide range of challenges and solutions, including data protection, database security, data compliance, and secure data transfer. Our identified solutions revolve around the use of encryption and secure container orchestration, complying with regulatory requirements such as GDPR, HIPAA, and secure data management strategies. Furthermore, identified solutions emphasizes the role of security measures such as access control and secure transmission protocols in ensuring data privacy, service continuity, fault tolerance, and the secure transfer of data between different cloud environments.
Access and Communication Control category emphasizes maintaining secure multi-cloud connectivity and managing user access control. The identified solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
Container Security category addresses issues like co-resident security, container isolation, and providing host security. Implementing secure placement strategies for containers across multi-cloud platforms, utilizing container isolation techniques, and enhancing manual defenses to mitigate risks of compromised containers are some of the identified solutions in this category.
Infrastructure and Deployment Management Security category emphasizes the significance of infrastructure and deployment management security. It encompasses aspects such as secure data deployment, infrastructure orchestration, storage security, and migration security. Essential solutions encompass the application of encryption techniques, implementation of access control mechanisms, adoption of standardized and secure infrastructure orchestration practices, and ensuring the secure migration of containerized applications across multi-cloud environment.
Security Monitoring and Breach Prevention category addresses a range of challenges and solutions related to enhancing security. This includes the implementation of automated security assessments, the use of vulnerability scanners, and the establishment of automated bootstrapping mechanisms for Virtual Machines (VMs) that incorporate security configurations and policies. Additionally, it involves deploying Security Information and Event Management (SIEM) systems, as well as Intrusion Detection System/Intrusion Prevention System (IDS/IPS) tools. The category also covers the application of log analysis solutions to proactively detect and prevent security breaches.
Trust and Compliance Management category identified solutions encompass on implementation of robust security controls, adherence to industry best practices and standards, and establishment of trust-based coalitions to foster secure community formation.
Placement Strategy: It recommends devising a security model for services hosted on volunteering fogs within multi-cloud environment, including formulation of secure communication protocols and implementation of access control mechanisms.
@@
Table 11. Security challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Data Security
S17 Data Protection Protecting data against threats in containerized multi-cloud setups. Multi-Cloud Data Protection through Secure Containers Deploys sensitive data in encrypted, isolated containers, using secure orchestration tools like Docker and Kubernetes.
S29 Database Security Securing databases within containerized multi-cloud environments. Enhanced multi-cloud Database Security via Containerization Secures databases by implementing encryption (e.g., Transparent Data Encryption) and robust access controls (e.g., RBAC, ABAC).
S37 Data Compliance Ensuring data processing and storage compliance across multi-cloud platforms. Multi-cloud Legal Compliance and Data Protection Maintains privacy and service quality by adhering to regulations (e.g., GDPR, HIPAA) and employing data protection strategies like encryption and audits.
S38 Data Security Adhering to data security standards in multi-cloud containerized systems. Secure CSP Interoperability and Selection in multi-cloud Tackles CSP compatibility issues and enhances security by implementing data handling practices and selecting compliant CSPs.
S53 Data Ownership and Privacy Tackles data ownership, privacy, and availability issues in multi-cloud setups. Multi-cloud Security via NoMISHAP Service Abstraction Uses NoMISHAP for service abstraction, enhancing privacy with encryption and pseudonymization, and ensuring availability with fault tolerance strategies like backups and failovers.
S58 Data Transfer Security Secures data during transfers in multi-cloud environments. Secure multi-cloud Integration via OpenStack Standardization Standardizes OpenStack APIs to prevent vendor lock-in, ensuring safe data transfers with encryption and secure protocols.
S85 Cloud Data Management Manages cloud data security, privacy, and compliance. Multi-cloud Data Security via DRA Framework Implements the DRA framework for comprehensive data management, ensuring encryption, access control, and compliance with standards.
Category 2: Access and Communication Control
S05 Connectivity and Access Control Ensures secure multi-cloud connectivity and robust user access control. Multi-cloud Containerized Application Security Implements TLS, JWT, OAuth 2.0 for secure authentication and authorization in multi-cloud setups.
S43 Secure Communication Secures communications across multi-cloud and MEC platforms. Middleware Secure Deployment for multi-cloud Uses middleware for secure cloud app deployment, with SSL/TLS and access control for secure communication.
S51 Provider-Agnostic Operations Facilitates secure, provider-agnostic operations in multi-cloud environments. Aneka-Based Security and Performance for Containerized multi-cloud Applications Employs Aneka for security and performance, offering transparent access and dynamic provider selection, while ensuring security via secure channels and monitoring.
Category 3: Container Security
S06 Co-resident Security Addresses security risks from container virtualization in multi-cloud environments. Secure Placement Strategy via Deep RL Uses deep reinforcement learning, like Kubernetes Cluster Autoscaler with RL, for secure container placement across multi-cloud platforms.
S25 Container Isolation Aims to prevent information leaks and power attacks through effective container isolation. Power and Leakage Management for multi-cloud Security Manages security via container isolation (e.g., Docker isolation) and Linux kernel-level resource isolation (e.g., SELinux).
S40 Container and Host Security Enhances security against compromised containers and strengthens defenses. Host-Container Cooperation for multi-cloud Defense Develops container-host defense strategies, utilizing tools like Falco, Sysdig, and practices for host hardening.
Category 4: Infrastructure and Deployment Management Security
S20 Data Deployment Secures private data deployment in multi-cloud environments. Secure Private Data Deployment Strategy Utilizes encryption, RBAC, and secure protocols (e.g., AES-256, SFTP, HTTPS) for safeguarding data deployment.
S44 Infrastructure Orchestration Orchestrates secure infrastructure for containerized applications. Standard Interface Implementation for Serverless Security Uses standard interfaces (e.g., OpenStack APIs) for secure, interoperable serverless application deployment across clouds.
S47 Storage Security Enhances security for multi-cloud storage in containerized setups. Multi-cloud Storage Security with MSSF Employs the MSSF for secure storage management, implementing RBAC, data encryption, and regular audits.
S80 Migration Security Ensures the secure migration of applications in multi-cloud environments. Infrastructure-Aware and Agnostic Secure Migration Leverages strategies and tools for safe migration, including image signing and data encryption, adhering to best practices like least privilege.
Category 5: Security Monitoring and Breach Prevention
S30 System Vulnerabilities Addresses vulnerabilities and patching delays in multi-cloud environments. Automated Security Evaluation and Bootstrapping Utilizes tools like OpenVAS and Nessus for security assessments, and automates VM setup with Terraform, Chef, Ansible, or Puppet.
S45 Breach Detection and Prevention Boosts breach detection and prevention in multi-cloud infrastructures. Metric Filtering and Fault Recovery for multi-cloud Security Enhances security with metric filtering, fault recovery, SIEM, IDS/IPS, and log analysis for proactive breach management.
Category 6: Trust and Compliance Management
S52 Application Security and Compliance Ensures security and compliance for applications across multi-clouds. Risk Analysis and Security SLAs with MUSA Framework Uses MUSA for risk analysis and security SLA management, incorporating secure coding, vulnerability management, and compliance audits.
S49 Data Security and Standardization Standardizes data security and cloud interfaces in multi-cloud environments. OpenStack API-based multi-cloud Security and Interoperability Utilizes OpenStack API for secure, standardized cloud interoperability, applying encryption, access control, and CIS benchmarks.
S57 Network Security Secures networks and trust in multi-cloud environments. Inter-cloud Communication Protection with SFC Employs Service Function Chaining (SFC) for secure inter-cloud communications, using encryption, integrity checks, and protocols like IPSec, WireGuard.
S83 Trust and Community Formation Builds secure communities and trust in multi-cloud settings. Trust-based Hedonic Coalitions for multi-cloud Security Forms trust-based coalitions for community security, leveraging IAM and trust frameworks for collaboration in multi-clouds.
S84 Orchestration and Adaptation Manages orchestration and adaptation for security and performance. Multi-cloud Service Orchestration and Cross-Level Adaptation Implements multi-cloud orchestration with tools like Kubernetes, Istio, and Prometheus for secure, adaptive application management.
Category 7: Placement Strategy
S33 Placement Strategy Secures service deployment on volunteer fogs in multi-cloud setups. Security Model for Volunteering Fog Services Creates a security model for fog-deployed services, incorporating SSL/TLS for communication and certificate or token-based access control.
@@ 3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
Table 12 presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment. The figure consists of eight categories, each focusing on different aspects of automation challenges and their solutions in a multi-cloud environment.
Multi-Cloud Automation: This category reports challenges and solutions related to automating processes for or combining and administering various cloud computing environments. It addresses challenges such as multi-cloud integration and provisioning, and provides solutions such as proxies for environment interface and model-driven engineering techniques.
Automation in Deployment and Scaling: This category reports challenges and solutions related to automating the deployment and scaling processes of containerized applications across multi-cloud platforms. It tackles challenges like container placement automation and application runtime switch automation. Solutions include intelligent container placement algorithms and service drivers for automated runtime switch.
Resource Management Automation: This category reports challenges and solutions related to automating the management of resources in multi-cloud environment. It addresses challenges such as resource overbooking management and containerized applications management. Solutions include ML-based resource overbooking detection and custom Kubernetes schedulers.
Data and Application Migration Automation: This category reports challenges and solutions related to automating the migration of data and applications in multi-cloud environment. It covers challenges such as data integrity during migration and legacy web applications migration. Solutions encompass the implementation of self-sufficient management cycles for streamlined deployment, alongside the utilization of adaptable architectural patterns designed for the modernization of legacy applications within a containerized multi-cloud environment.
Testing and Benchmarking Automation: This category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment. It addresses challenges like user-oriented testing automation and benchmarking automation. Solutions include container-based lightweight virtualization for testing and smart Docker benchmarking orchestrators.
Standardization and Interoperability Challenges: This category reports challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments. It covers challenges like standardization in automation processes and cloud provider selection. Solutions include service mesh networks and AI-driven cloud provider selection tools.
Application and Service Management: This category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment. It addresses challenges such as 5G ecosystem component selection and end-to-end tail latency management. Solutions include open-source software components for customizable orchestration and modeling approaches for performance evaluation.
Runtime and Service Discovery: This category deals with challenges and solutions related to runtime management and service discovery in multi-cloud environment, which involves managing the complexity of network control and administration within containerized applications across multiple cloud environments.
@@
Table 12. Automation challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
@@
S10 Multi-Cloud Integration Automating processes to manage the complexities of integrating multiple cloud environments Proxies for Environment Interface Proxies for individual environment interface, unified view of resources, shielding from underlying complexity in multi-cloud environment
S20 Multi-Cloud Provisioning and Adaptation Automating provisioning, deployment, and dynamic adaptation of multi-cloud applications for efficiency and timely response Model-driven Engineering Techniques Model-driven engineering techniques for dynamic provisioning, deployment, and adaptation in multi-cloud environment
S33 Fog Creation in Multi-Cloud Automating the dynamic creation of nearby fog in multi-cloud environment Dynamic On-demand Fog Computing Dynamic on-demand fog computing architecture in multi-cloud environment
S47 Intelligent User-Friendly Automation Addressing the lack of intelligent, user-friendly automation in delivery platforms MSSF MSSF (Multi-cloud Storage Selection Framework) for automatic selection of multi-cloud redundancy storage based on user needs and security rules
S44 Multi-Stack Orchestration Automating the specification, coordination, and orchestration of different infrastructure stacks and cloud services TOSCA-based Serverless Deployment Topology and Orchestration Specification for Cloud Applications (TOSCA) for automated deployment of multi-cloud serverless applications
S34 Secure Multi-Cloud App Creation Providing tools and mechanisms for non-security expert developers to automate secure multi-cloud app creation MUSA Framework MUSA framework for application modeling, risk analysis, cloud service selection, composite SLA generation in multi-cloud environment
S37 Unified Abstraction for Multi-Cloud Deployment Implementing a common abstraction for products offered in automated multi-cloud deployments Application-level Resource Managers Application-level resource managers, native multi-cloud support for enterprise applications in multi-cloud environment
S52 Multi-Cloud Deployment Integration Addressing integration complexity in multi-cloud deployment automation MUSA Deployer Tool MUSA Deployer tool for automation of deployment process, deployment plan, and security compliance in multi-cloud environment
S57 Deployment Across Federated Clouds Automating deployment across federated cloud providers Orchestrated Deployment Subsystem Orchestrated deployment subsystem, Beacon Broker, and federated environments for coordinated application deployment in multi-cloud environment
S58 Hybrid/Multi-Cloud Setup Automating the setup of hybrid/multi-cloud environment to address security, vendor lock-in, and interoperability issues Interoperability Standardization Interoperability standardization and federation of clouds in multi-cloud environment
S80 Multi-cloud Interoperability Facilitating multi-cloud interoperability, application topology definition/composition, and elastic runtime adaptation for containerized services Multi-cloud Interoperability Multi-cloud interoperability, application topology definition/composition, and elastic runtime adaption in multi-cloud environment
S81 Multi-cloud Brokerage System Overcoming API heterogeneity and promoting semantic interoperability for a brokerage system in multi-cloud container management. Managing Large Number of Cloud Services Interoperability and vendor lock-in caused by the diversity of cloud solutions, complexity of multi-cloud systems administration
S84 Multi-cloud Orchestration Implementing multi-cloud business process provisioning and cross-level orchestration for cloud services in a containerized setting Manual Migration to the Cloud Manual migration of business processes to the Cloud, limited multi-cloud and cross-cloud orchestration
S85 DevOps for Multi-cloud Building a DevOps reference architecture with a CI broker for automation within a containerized multi-cloud environment Agile Software Development Automation Automation of agile software development, application deployment using CI-broker, automated faster application delivery, automated and integrated testing of the application in multi-cloud environment

### v2

@@ 3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
This section presents a framework for addressing security challenges in containerized applications in multi-cloud environment. Based on a detailed review, it identifies seven key categories of challenges and solutions, outlined in Table LABEL:tab:Security-challenges-solutions. Each category is briefly reported below:
Data Security in Container-based Applications category covers multiple challenges related to data security in container-based applications, including data protection, database security, data compliance, and secure data transfer in a multi-cloud environment. Out of the 121 reviewed studies, 11 specifically highlighted these issues and provided corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions focus on implementing robust encryption techniques, secure container orchestration, and compliance with regulatory standards such as GDPR and HIPAA. The solutions also include using Role-Based and Attribute-Based Access Controls (RBAC and ABAC) to manage secure access effectively. Standardization efforts like Secure CSP Interoperability and OpenStack API integration aim to mitigate risks of unauthorized access and vendor lock-in. Furthermore, comprehensive frameworks like DRA emphasize secure data management, ensuring privacy, fault tolerance, and secure data transfers across different cloud platforms.
Table 11. Identified security challenges and solutions for containerized applications in multi-cloud environments
ID Challenge Solution
Category 1: Data Security
(Slominski et al., 2015) Data Protection Multi-Cloud Data Protection through Secure Containers
(Gupta et al., 2021) Database Security Enhanced multi-cloud Database Security via Containerization
(Dimitrijevic et al., 2019) Data Compliance Multi-cloud Legal Compliance and Data Protection
(Pandey et al., 2022) Data Security Secure CSP Interoperability and Selection in multi-cloud
(Acquaviva et al., 2017) Data Ownership and Privacy Multi-cloud Security via NoMISHAP Service Abstraction
(Sitaram et al., 2018) Data Transfer Security Secure multi-cloud Integration via OpenStack Standardization
(Bou Ghantous and Gill, 2021) Cloud Data Management Multi-cloud Data Security via DRA Framework
(Ahmad and Aujla, 2023) Data leaks and access risks Cryptography, smart contracts, encryption
(Sofia et al., 2024) Decentralized data security challenge Secure data handling with decentralized AI
(Neeraj et al., 2023) SLA data integrity challenge Blockchain for SLA security and integrity
(Santos et al., 2023) Multi-cloud data privacy issue Stringent security controls for data protection
Category 2: Access and Communication Control
(Gattobigio et al., 2022) Connectivity and Access Control Multi-cloud Containerized Application Security
(Dreibholz et al., 2019) Secure Communication Middleware Secure Deployment for multi-cloud
(Buyya and Barreto, 2015) Provider-Agnostic Operations Aneka-Based Security and Performance for Containerized multi-cloud Applications
(Hossain et al., 2023) Microservices privacy and access control Encryption, authentication, service mesh for security
(Fugkeaw, 2023) Decentralized IAM security issue Blockchain for IAM security and data integrity
(Roy and Kalafatis, 2023) Secure communication across edge-cloud environments WireGuard for secure edge-cloud communication
(Benmerar et al., 2023) Multi-domain edge communication security Secure communication and AI anomaly detection
(Huang and Wang, 2023) Secure robot-cloud communication Secure communication and encryption for data integrity
Category 3: Container Security
(Deng et al., 2022) Co-resident Security Secure Placement Strategy via Deep RL
(Gao et al., 2017) Container Isolation Power and Leakage Management for multi-cloud Security
(Bélair et al., 2019) Container and Host Security Host-Container Cooperation for multi-cloud Defense
(Wang et al., 2023) Container security and malware detection DockerWatch for malicious activity detection
(Tiwari and Sharma, 2023) Multi-tenant serverless security challenge Security best practices for serverless deployments
Category 4: Infrastructure and Deployment Management Security
(Ferry et al., 2014) Data Deployment Secure Private Data Deployment Strategy
(Wurster et al., 2018) Infrastructure Orchestration Standard Interface Implementation for Serverless Security
(Libardi et al., 2014) Storage Security Multi-cloud Storage Security with MSSF
(Quint and Kratzke, 2018) Migration Security Infrastructure-Aware and Agnostic Secure Migration
(Solayman and Qasha, 2023) Distributed infrastructure security challenge Orchestration security for IoT data protection
(Kataru et al., 2023) Securing AWS app deployments AWS best practices for CI/CD security
Category 5: Security Monitoring and Breach Prevention
(Sheridan et al., 2017) System Vulnerabilities Automated Security Evaluation and Bootstrapping
(Trihinas et al., 2015) Breach Detection and Prevention Metric Filtering and Fault Recovery for multi-cloud Security
(Lee and Nam, 2023) Network policy misconfiguration risks Automated network policy discovery and verification
(Nguyen and Chien, 2023) Stream processing security risks RBAM framework for secure data handling
Category 6: Trust and Compliance Management
(Casola et al., 2017) Application Security and Compliance Risk Analysis and Security SLAs with MUSA Framework
(Harwalkar et al., 2019b) Data Security and Standardization OpenStack API-based multi-cloud Security and Interoperability
(Tricomi et al., 2017) Network Security Inter-cloud Communication Protection with SFC
(Wahab et al., 2016) Trust and Community Formation Trust-based Hedonic Coalitions for multi-cloud Security
(Kritikos et al., 2019) Orchestration and Adaptation Multi-cloud Service Orchestration and Cross-Level Adaptation
(Petrovski and Gusev, 2024) Multicloud compliance and security issue AWS security measures for multicloud environments
(Alonso et al., 2023a) Cross-cloud security and compliance Security frameworks for multi-cloud environment
Category 7: Placement Strategy
(Sami and Mourad, 2020) Placement Strategy Security Model for Volunteering Fog Services
(Pandi et al., 2023) Multi-cloud security concerns Security groups, VPCs, and infrastructure security
(Ramesh et al., 2023) Multicloud environment security issue Multicloud security for risk diversification
(Sofia et al., 2023) Edge-cloud security concerns Enhanced security protocols for decentralized environments
(Harzenetter et al., 2023) Cross-technology update security risks Integrated security in workflows for management
(Nagarajan et al., 2023) Resource allocation security risks Security measures in resource allocation framework
Access and Communication Control category reports on multiple challenges related to maintaining secure multi-cloud connectivity and managing user access control. Out of the 121 reviewed studies, 9 specifically identified issues and proposed solutions addressing secure communication, provider-agnostic operations, and decentralized identity management (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). Advanced encryption and authentication mechanisms, e.g., TLS, JWT, and OAuth 2.0, were identified as solutions to secure communications across cloud and edge environments. Secure deployment of applications using middleware solutions was also discussed, which make use of SSL/TLS and access controls to enhance security attributes. Provider-agnostic middleware, e.g., the one based on Aneka, enables runtime provider selection and thus facilitates seamless operations across multi-cloud environment. Solutions for decentralized IAM security relate to blockchain utilization for improving data integrity and confidentiality. Secure communications in multi-domain edge environments are achieved through frameworks like WireGuard, backed by strong service mesh systems that maintain connectivity and protect against unauthorized access. Security monitoring for anomaly detection and policy enforcement was also noted as important for sustaining secure operations.
Container Security category reports on challenges such as co-resident security, container isolation, host security, and multi-tenant serverless security in multi-cloud environment. Among the 121 reviewed studies, 5 identified these issues and proposed solutions to address them (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The solutions identified include the use of secure placement strategies via deep reinforcement learning to manage secure container placement in multi-cloud platforms. Security measures for container isolation include the technique of Power and Leakage Management, followed by Linux kernel-level isolation using SELinux. In order to enhance security on both container and host sides, solutions such as Host-Container Cooperation and Docker-based detection tools, like DockerWatch, enhance the defenses and malicious activity detection. Finally, with a view to further securing multi-tenant serverless environments, the implementation of best practices regarding security in the CLI tool used for managing serverless deployments was underlined.
Infrastructure and Deployment Management Security category reports on challenges related to securing data deployment, infrastructure orchestration, storage security, and migration security in multi-cloud environment. Out of the reviewed 121 studies, 6 specifically addressed these types of challenges and proposed solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions include the use of a Secure Private Data Deployment Strategy with encryption techniques like AES-256 and secure protocols such as SFTP and HTTPS to safeguard data. In the case of infrastructure orchestration, the focus is on the implementation of standard interfaces, such as OpenStack APIs, to ensure security and interoperability during deployments. For storage security, the solutions introduced the Multi-Cloud Storage Framework (MSSF), providing secure storage management through the implementation of RBAC and data encryption, along with periodic audits. The solutions also introduced measures for secure migration through both Infrastructure-Aware and Agnostic Secure Migration strategies, emphasizing image signing and data encryption under the least-privilege principle. Additionally, the reviewed studies proposed orchestration security for IoT distributed data and operations, as well as AWS best practices to help protect CI/CD pipelines running in multi-cloud environment.
Security Monitoring and Breach Prevention category reports on challenges related to system vulnerabilities, breach detection and prevention, and network policy misconfiguration risks in multi-cloud environment. Out of the 121 reviewed studies, 4 specifically addressed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions suggest implementing automated security evaluations and bootstrapping mechanisms for Virtual Machines with security configurations using tools like Terraform, Chef, Ansible, or Puppet. The studies also highlighted several other tools for vulnerability assessment, such as OpenVAS and Nessus, which can be used to address delays in patching and reduce system vulnerabilities. In terms of breach detection and prevention, the reviewed studies emphasized the need to enhance security through metric filtering, fault recovery, and the automated deployment of SIEM, IDS, and IPS tools. Log analysis solutions were considered necessary for proactive breach management. As a result, automated network policy discovery and verification mechanisms, such as KUNERVA, were proposed, incorporating rigorous verification to prevent most potential breaches in network policy management.
Trust and Compliance Management category reports on solutions related to ensuring security and compliance in multi-cloud environment. Among the 121 reviewed studies, 7 specifically discussed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The identified solutions emphasize the need for security controls and standardized practices, such as risk analysis and secure coding using MUSA. The use of standardized API-based cloud interfaces, like the OpenStack API, was also emphasized to ensure both data security and interoperability between different cloud platforms. Establishing trust-based coalitions for community security was another important approach, supported by IAM systems and secure communication frameworks like SFC to enable secure interactions and collaborations within multi-cloud settings.
Placement Strategy category highlights the need for secure service deployment on volunteering fog nodes in multi-cloud environment. Of the 121 reviewed studies, 6 proposed solutions for these challenges (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The solutions involve developing a security model that integrates SSL/TLS protocols for communication and certificate-based access controls to protect deployed services on fog nodes. The strategies include implementing a combination of security groups, VPCs, and infrastructure-level measures to distribute risks and prevent unauthorized access. The reviewed studies recommend adaptive methods for handling security in dynamic and decentralized edge-cloud environments by proposing cross-technology security updates and resource allocation frameworks to avoid conflicts and vulnerabilities.
@@ 3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
Table LABEL:tab:Automation-solutions presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment. The figure consists of eight categories, each focusing on different aspects of automation challenges and their solutions in a multi-cloud environment.
Multi-Cloud Automation category reports on challenges and solutions related to automating processes for combining and administering various cloud computing environments. Out of the 121 reviewed studies, 13 identified these challenges and provided associated solutions (see Table LABEL:tab:Automation-solutions and Automation Challenges-Solutions sheet in (Waseem et al., 2024)). Key challenges involve the integration, provisioning, orchestration, and adaptation of multiple clouds. The solutions focus on utilizing proxies for environment interfaces to ease management across multiple cloud platforms and employing model-driven engineering techniques to manage provisioning and adaptation more efficiently. Dynamic on-demand fog computing was identified as an essential approach for handling multi-cloud environment dynamically. At the level of cloud redundancy and resource management driven by security policies, smart and user-friendly automation techniques, such as MSSF, were recommended. For managing the complexity of multi-cloud orchestration, serverless deployment methodologies were proposed, using models like TOSCA-based orchestration to improve scalability and coordination between different cloud stacks. Recommendations also include deploying application-level resource managers to support multi-cloud operations and adopting secure abstraction models to create uniform interfaces for managing a wide range of cloud products. MUSA Deployer tools were highlighted for their key role in automating deployment and monitoring activities, ensuring overall security compliance, and simplifying the deployment pipeline across federated clouds.
Automation in Deployment and Scaling category reports on challenges and solutions related to automating deployment and scaling processes for containerized applications across multi-cloud platforms. Out of the 121 reviewed studies, 12 identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). Key challenges include intelligent container placement, application runtime switch automation, and service deployment issues. The proposed solutions emphasized intelligent algorithms for container placement, using machine learning for multi-cloud workload balancing. Runtime switch techniques are also being automated with methods such as service drivers, which are designed to be lightweight and enable seamless application migration and replication. DevOps methodologies were frequently mentioned in the literature in relation to managing deployment and scaling through continuous development and error reduction. Security during deployment is ensured through automated security measures integrated into the pipeline, aiming for a secure setup with reduced vulnerabilities. The automation strategies include elastic resource provisioning using predictive analytics to handle fluctuating workloads in real time. Optimization models were involved in solutions for virtual function placement and container deployment to enhance efficiency. This emphasis on automating and optimizing deployment strategies led to the proposal of comprehensive multi-cloud orchestration tools that facilitate the deployment of containers and VMs, aimed at improving agility and scalability within diverse cloud environments.
@@
Table 12. Identified automation challenge and solution for containerized applications in multi-cloud environment
ID Challenge Solution
@@
(Sabbioni et al., 2020) Integrating multi-cloud environment Implementing Environment Proxies
(Ferry et al., 2014) Automating Multi-Cloud Provisioning Adopting Model-Driven Engineering
(Sami and Mourad, 2020) Creating On-Demand Fog in Multi-Cloud Enabling On-Demand Fog Computing
(Libardi et al., 2014) Enhancing User-Friendly Automation Using Multi-Cloud Storage Selection
(Wurster et al., 2018) Orchestrating Multi-Stack Environments Deploying with TOSCA-based Approach
(Rios et al., 2017) Securing Multi-Cloud Application Creation Applying MUSA Framework
(Dimitrijevic et al., 2019) Unifying Resource Abstraction Managing with Resource Managers
(Casola et al., 2017) Integrating Multi-Cloud Deployments Automating MUSA Deployment
(Tricomi et al., 2017) Coordinating Federated Cloud Deployments Orchestrating Subsystem Deployment
(Sitaram et al., 2018) Setting Up Hybrid multi-cloud environment Standardizing Interoperability
(Quint and Kratzke, 2018) Ensuring Multi-Cloud Interoperability Enabling Multi-Cloud Support
(Challita et al., 2017) Managing Multi-Cloud Brokerage Systems Managing Cloud Services
(Kritikos et al., 2019) Implementing Multi-Cloud Orchestration Migrating Manually to Cloud
(Bou Ghantous and Gill, 2021) Establishing DevOps in Multi-Cloud Automating Agile Development
(da Silva et al., 2023) Addressing Cost and Migration Challenges Managing Complexity
(Pandi et al., 2023) Overcoming Multi-Cloud Automation Barriers Integrating Jenkins and Kubernetes
(Ramesh et al., 2023) Improving Multi-Cloud Cost Efficiency Optimizing Deployment Costs
(Petrovski and Gusev, 2024) Simplifying Multi-Cloud Deployment Complexity Adopting AWS Deployment Strategies
(Kataru et al., 2023) Maintaining Multi-Cloud Applications Automating Resource Allocation